Four point charges are equally separated by 1.0 m in air with q1= 30 nC, q2 = 4.0 nC, q3 = 2.0 nC, q4 = 5.0 nC. (a) What is the potential at a point halfway between q1 and q2? (b) How much work is needed to bring a 6 nC charge from infinity to this point?

1
Expert's answer
2022-03-15T10:36:42-0400

a)

"\varphi=\frac{k(q_2-q_1)}{\frac r2}=\frac{2k(q_2-q_1)}r=18~V,"

b)

"A=q\varphi =0 .1~\mu J."